OpenStudy (anonymous):

true or false and why please help! 2+3 not equal to 5 if and only 10-1= 3^2

OpenStudy (anonymous):

thats false because 2 plus 3 does equal five and 10-1 does equal 3 squared...

OpenStudy (kainui):

What is the confusing part of this question?

OpenStudy (anonymous):

is it that simple?

OpenStudy (kainui):

Yep! You can rephrase it in your mind as: A true statement (3+2=5) is not true if and only if another true statement is true (10-1=3^2) Try to think of it in terms of that, and don't overthink them! =D

OpenStudy (anonymous):

wait so how should i explain it @Kainui

OpenStudy (anonymous):

the answer is false?

OpenStudy (kainui):

Yes it's false because 2+3=5 is a completely independent of 10-1=3^2

OpenStudy (anonymous):

what do you mean completely independent @Kainui

OpenStudy (anonymous):

take a look at the truth table for iff statements http://en.wikipedia.org/wiki/If_and_only_if